# Characterizing the Sensitivity to Individual Bit Flips in Client-Side Operations of the CKKS Scheme

### v1

@@ IV-G Evaluation Metrics
The element-wise error is defined as:
$\delta_{i}=y_{i}-x_{i}$
and the MSE is computed as:
$\mathrm{MSE}=\sqrt{\frac{1}{k}\sum_{i=1}^{k}\delta_{i}^{2}}$
@@
For each decrypted element, we compute the relative error:
$\epsilon_{i}=\frac{|y_{i}-x_{i}|}{|x_{i}|}$
Results are classified into four categories based on the relative error $\epsilon_{i}$, with each category corresponding to a different tolerance level, which may vary depending on the intended application of FHE:
  1. 1.
    Severe Error ($\epsilon_{i}>10$): the result differs by more than one order of magnitude from the correct value, often rendering it unusable.
  2. 2.
    Significant Error ($0.1<\epsilon_{i}\leq 10$): the result may be unacceptable for precision tasks, but tolerable in coarse-grained applications such as visual representations or robust classification.
  3. 3.
    Moderate Error ($0.01<\epsilon_{i}\leq 0.1$): the deviation is small and potentially acceptable in tasks like statistical analysis or machine learning where high fidelity is not required.
  4. 4.
    Low Error ($\epsilon_{i}\leq 0.01$): the result is nearly indistinguishable from the correct value and is considered acceptably decrypted for most practical purposes.
V Experimental Results
This section presents the results obtained by injecting individual bit-flips during encryption in the CKKS scheme, as described in Section IV, with the goal of evaluating the scheme’s resilience against hardware- or software-induced faults.
Experimental campaigns were carried out using both HEAAN 1.0 and OpenFHE 1.2.3, in their vanilla and optimized configurations (with RNS and NTT).
V-A Vanilla CKKS
V-A1 Evolution of the $L_{2}$ norm under Single Bit‑Flips
To deepen the comparison, we repeated the experiments with HEAAN 1.0, obtaining results that match our theoretical predictions closely. This direct comparison between HEAAN and OpenFHE revealed a marked difference between the vanilla implementation and the fully optimized variant with RNS and NTT: the latter exhibits noticeably greater sensitivity to bit-flips, as evidenced by higher $L_{2}$ error magnitudes across most bit positions.. To isolate the effect of these optimizations, we therefore configured OpenFHE in a simulated vanilla mode by disabling both NTT and RNS, as described in Sec. IV-B.
In order to validate this simulation, we ran both HEAAN and OpenFHE (vanilla) under identical conditions. Because OpenFHE only accepts real inputs, we also restricted HEAAN accordingly to real inputs and, after decryption, considered only the real part of each result. This ensured a homogeneous comparison between the two libraries and confirmed that OpenFHE (vanilla) faithfully reproduces the behavior of HEAAN 1.0, in both trend and magnitude.
Figure 7 shows the case of OpenFHE simulated as CKKS vanilla.
Figure 7: Impact of a single-bit flip during encryption in OpenFHE without RNS or NTT. The horizontal axis indicates the position of the flipped bit within the ciphertext, while the vertical axis shows the resulting $L_{2}$ norm of the decoding error. Parameters: 60‑bit modulus, scaling factor $\Delta=2^{50}$, and ring dimension $N=4$.
We observe that, for the first polynomial $c_{0}$, the behavior mirrors the theoretical analysis in Fig. 4: the least significant bits exhibit higher robustness against bit‑flips, followed by an exponential increase in error magnitude.
In the case of OpenFHE, the expected absolute robustness at coefficient $N/2$ does not hold. During decoding, the coefficient vector prior to the DFT (a specialized FFT) is injected with Gaussian noise affecting both its real and imaginary components. This noise injection is a security measure to thwart key‑recovery attacks.
The introduced Gaussian noise has a standard deviation corresponding to the dispersion of the imaginary part of the vector being transformed, which—by properties of the scheme—also follows a Gaussian distribution. Flipping any bit of any coefficient, including the $N/2$ coefficient, increases this noise standard deviation, especially when higher‑order bits are modified, leading to increased noise variance that propagates across all elements of the vector, significantly altering the final output and negating the robustness at $N/2$.
We experimentally confirmed that, when this noise injection is disabled, the observed behavior aligns with theoretical expectations: the coefficient $N/2$ of $c_{0}$ becomes robust to bit‑flips. Conversely, since HEAAN 1.0 does not perform this noise injection, its sensitivity matches the theory exactly.“‘
V-A2 Impact of the Scaling Factor on Error Categorization
One of the fundamental parameters of the CKKS scheme is the scaling factor $\Delta$. As analyzed in Sec. III, increasing $\Delta$ reduces the relative impact of a single‐bit flip during encoding or encryption and also increases the proportion of bits that remain resilient to such faults.
To investigate this effect in practice, we repeated the previous single‐bit‐flip analysis using four different scaling factors, namely $\Delta\in\{2^{20},2^{30},2^{40},2^{50}\}$. Figure 8 shows the $L_{2}$ norm of the decoding error for a bit flip in the first ciphertext polynomial $c_{0}$ when running OpenFHE in simulated vanilla mode (i.e., with RNS and NTT disabled). For clarity, only results for $c_{0}$ are shown, since the behavior for $c_{1}$ is analogous.
Figure 8: Impact of a single‐bit flip during encryption in OpenFHE without RNS or NTT. The x‑axis indicates the bit position in ciphertext $c_{0}$, while the y‑axis shows the resulting $L_{2}$ norm of the decoding error on a symmetric log scale. Parameters: 60‑bit modulus, scaling factors $\Delta=2^{20},2^{30},2^{40},2^{50}$, and ring dimension $N=8$.
We observe that higher values of $\Delta$ yield lower $L_{2}$ norms after a bit flip, indicating stronger attenuation of induced errors. Furthermore, the point at which the error norm begins its exponential growth shifts in alignment with the magnitude of $\Delta$, confirming that larger scaling factors increase the number of error-resilient bit in each coefficient’s representation.
@@ V-A2 Impact of the Scaling Factor on Error Categorization
For this experiment, we adopted a realistic setting with ring dimension $N=2^{13}=8192$ (i.e., 4096 encrypted slots). For each $\Delta\in\{2^{20},2^{30},2^{40},2^{50}\}$ and each single‐bit flip, we counted how many slots were correctly decrypted and assigned the result to one of the three categories. Figure 9 presents a histogram of the category frequencies for each scaling factor.
@@
[…]
V-A3 Effect of the Effective Representation Range on Error Categorization
We now analyze the impact of the effective dynamic range of the input data representation. We fix $\Delta=2^{20}$, which allows for a wider input range. Recall that the number of bits available for the integer part is approximately given by $q_{0}-\Delta$. The effective dynamic range refers to the ratio between the largest and smallest values that can be represented in fixed-point encoding. In our experiments, this range is controlled indirectly by limiting the number of bits allocated to the fractional part of the input values.
[…]
We observe behavior similar to the impact of the scaling factor, as seen in Fig. 9. As we increase the size of the input space, the percentage of correctly decrypted elements increases while the percentage of incorrectly decrypted elements decreases. If the size of input elements grows while maintaining a fixed scaling factor, this allows representing values with greater relative precision compared to the existing noise.
V-A4 Effect of reducing the number of encrypted elements
As noted previously, CKKS can encode up to $N/2$ slots, where $N$ is the ring degree (a power of two). When we encode fewer slots—still a power of two—we observe that additional coefficients remain unaffected by single‐bit flips. Figure 11 shows results for $N=16$ with only $N/4=4$ encrypted slots, allowing us to inspect a larger set of coefficients and evaluate their sensitivity.
Figure 11: Impact of a single‐bit flip during encryption in OpenFHE without RNS or NTT. The horizontal axis indicates the bit position in the ciphertext, while the vertical axis shows the resulting $L_{2}$ norm of the decoding error. Parameters: 60‑bit modulus, scaling factor $\Delta=2^{20}$, ring dimension $N=16$, and 4 encrypted slots.
The results reveal that odd‐indexed coefficients of polynomial $c_{0}$ remain unchanged under bit‐flips, whereas even‐indexed coefficients exhibit the same sensitivity patterns observed earlier.
This phenomenon stems from CKKS’s decoding process in both OpenFHE and HEAAN. As explained in Sec. II, the ratio between the maximum number of slots ($N/2$) and the number actually used defines an interval—termed the gap—between the coefficients participating in the inverse transform. During the “special” FFT, only coefficients whose indices are multiples of this gap are considered and scaled in the input vector (see Sec. I for details).
Consequently, when encoding $N/2$ slots, all $N$ coefficients of $c_{0}$ are sensitive to bit‐flips; encoding $N/4$ slots reduces the number of sensitive coefficients of $c_{0}$ to $N/2$; and so on. In the case of $c_{1}$, no matter the number of slots used, all coefficients are sensitive to bit changes. Experimental campaigns confirm that our OpenFHE vanilla simulation (with RNS and NTT disabled) matches the expected behavior.
V-B Resilience under an intermediate RNS mode without NTT
We next conducted experimental campaigns to assess CKKS’s sensitivity when using only one of the numeric optimizations—either NTT or RNS. I n both pure-NTT and pure-RNS configurations, any single‐bit flip caused decryption errors of unacceptable magnitude. As shown in our theoretical analysis, bit flips under RNS produce errors several orders of magnitude larger than under NTT, while NTT alone still yields extremely large errors that preclude practical precision.
An interesting case emerges when CKKS is run with RNS only and the number of encrypted slots is halved or more. Figure 12 reports results for ring dimension $N=16$ with only 444 encrypted slots (i.e., one power‐of‐two below the maximum), scaling factor $\Delta=2^{20}$, and 8‐bit input values.
[…]
Polynomial $c_{1}$ exhibits no robust bits at all: every bit flip causes uniform decoding failure, so its error‐categorization histogram is flat and standard deviation is zero.
These results demonstrate that running CKKS with RNS only—together with a non‐maximal but non‐zero gap—can concentrate robustness on a subset of coefficients. Such a configuration suggests a trade‐off between ring utilization and fault resilience. Future work may explore tailored RNS and NTT parameterizations to minimize sensitivity to bit flips while preserving sufficient slot capacity, laying the groundwork for error‐reduction strategies in homomorphic encryption schemes that balance precision and fault tolerance.“‘
V-C Summary and Comparison
We provides a high‑level comparison of the different CKKS configurations evaluated in this work. For each scenario, we highlight the key resilience pattern and any notable observations.
  • Vanilla CKKS:
    • Least‐significant bits exhibit robustness; error grows exponentially for higher bits.
    • Coefficient $N/2$ robust in HEAAN 1.0 but not in OpenFHE vanilla (due to Gaussian noise injection).
  • Varying $\Delta$:
    • Larger $\Delta$ delays the onset of exponential error growth to more significant bits.
    • $\Delta\geq 2^{30}$ yields $\mathbf{Tail}$ outcomes in over 80% of trials.
  • Effective Representation Range:
    • Increasing integer‐part range raises $\mathbf{Tail}$ frequency and reduces $\mathbf{Head}$ frequency.
    • Behavior parallels that of increasing $\Delta$.
  • Reduced Slots:
    • Encoding fewer slots reduce the number of sensitive coefficients in $c_{0}$.
  • RNS without NTT:
    • Some coefficients of $c_{0}$ remain robust; all flips in $c_{1}$ fail.
    • Demonstrates a “gap”‐driven concentration of robustness when using RNS only.
This concise overview highlights how each parameter adjustment—from vanilla through optimized variants—modulates CKKS’s fault tolerance, guiding practical parameter selection under two‐column constraints.

### v3

@@ IV-E Evaluation Metrics
We define the element-wise error as $\delta_{i}=y_{i}-x_{i}$, and the MSE is computed as:
$\mathrm{MSE}=\sqrt{\frac{1}{k}\sum_{i=1}^{k}\delta_{i}^{2}}$ (18)
@@
For each decrypted element $i$, we compute the relative error:
$\epsilon_{i}=\frac{|y_{i}-x_{i}|}{|x_{i}|}=\frac{|\delta_{i}|}{|x_{i}|}$ (19)
To determine whether a vector is correctly recovered, we use the metric in Eq. (19) and proceed element-wise. For each element $i$, we compute the relative error with the reference value (i.e, same element of the vector recovered wihout no bit flip). We then classify the outcome of the procedure according to the fraction of elements correctly recovered:
@@
IV-F Experimental Setup
All experiments are execute on a dedicated workstation equipped with an Intel Core i7-11700 CPU (8 physical cores, 16 threads, 16 MB L3 cache, base 2.50 GHz) and 32 GB DDR4 RAM (dual channel, 3200 MHz). The operating system is Arch Linux (kernel Linux 6.9.7-arch1-1, distribution snapshot 257.5-1). We compiled both OpenFHE 1.3.0 and HEAAN 1.0 in release mode using g++ 15.2.1 with flags -O3. For HEAAN, NTL (NTL 11.5.1) was linked dynamically; OpenFHE was built with its default backend (no GPU acceleration).
V Experimental Results
This section presents the results obtained from injecting single bit flips during the encryption stage of the CKKS scheme, as described in Section IV, in order to evaluate its resilience against induced faults.
V-A Vanilla CKKS
We first analyze the vanilla CKKS scheme using OpenFHE; i.e., without RNS and NTT optimizations.
V-A1 Evolution of the $L_{2}$-norm Under Single Bit Flips
To evaluate the fault sensitivity of the vanilla CKKS variant, we systematically flipped one bit at a time in the ciphertexts generated during encryption, covering all bit positions in both $c_{0}$ and $c_{1}$ polynomials. After each injection, the modified ciphertext was decrypted and decoded, and the $L_{2}$-norm of the difference with respect to the fault-free output was computed. This metric quantifies the magnitude of the decoding error associated with each bit position.
Figure 8 shows the resulting $L_{2}$-norm distribution for all possible single-bit flips, highlighting which regions of the ciphertext are most sensitive to faults. The ring dimension is $N=4$, with modulus $q_{0}=60$-bit and scaling factor of $\Delta=2^{50}$.
Figure 8: Bit flip during encryption on vanilla mode. The horizontal axis indicates the position of the flipped bit within the ciphertext, while the vertical axis shows the resulting $L_{2}$-norm of the decoding error. Parameters: 60‑bit $q_{0}$, input range from 0 to 2560256, $\Delta=2^{50}$, and $N=4$.
We observe that, for the first polynomial $c_{0}$, the behavior closely follows the theoretical trend shown in Figure 4: least significant bits exhibit higher resilience to bit flips, while the error magnitude grows exponentially from the bit position corresponding to $\Delta$ up to the bit position defined by the modulus $q_{0}$, after which it abruptly drops to zero.
In our theoretical analysis, as well as in HEAAN, the coefficient $N/2$ of $c_{0}$ is completely robust. However, in OpenFHE, as we can observe in Figure 8 near bit 192, this behavior does not occur. This difference arises because OpenFHE, during decoding, injects Gaussian noise into the plaintext prior to the DFT (a specialized FFT), perturbing both real and imaginary components as a countermeasure against key-recovery attacks.
The injected noise has a standard deviation proportional to the spread of the imaginary part of the pre-DFT vector (plaintext), which, due to the scheme’s properties, follows a Gaussian distribution. Flipping any bit in any coefficient, including $N/2$, alters this spread; higher-order bit flips have a stronger effect, increasing the overall noise variance and propagating it across all polynomial coefficients, thereby removing the theoretical robustness of coefficient $N/2$. Disabling this noise injection in OpenFHE restores the expected behavior: coefficient $N/2$ of $c_{0}$ becomes robust to bit-flips. Since HEAAN 1.0 does not include this security noise, its experimental sensitivity matches the theoretical prediction exactly.
V-A2 Impact of the Scaling Factor
One of the key parameters of the CKKS scheme is the scaling factor $\Delta$. As analyzed in Section III, increasing $\Delta$ reduces the relative impact of a single‐bit flip during encoding or encryption and also increases the proportion of bits that remain resilient.
To validate this behavior experimentally, we analyze the impact of three different scaling factors, namely $\Delta\in\{2^{20},2^{40},2^{50}\}$. Figure 9 shows the $L_{2}$-norm of the decoding error for a bit flip in the first ciphertext polynomial $c_{0}$ on vanilla mode. For clarity, only results for $c_{0}$ are shown, since the behavior in $c_{1}$ is analogous (as is shown in Fig. 8).
Figure 9: Bit flip during encryption on vanilla mode. The horizontal axis indicates the bit position in ciphertext $c_{0}$, while the vertical axis shows the resulting $L_{2}$-norm of the decoding error on a symmetric log scale. Parameters: 60‑bit $q_{0}$, $\Delta\in\{2^{20},2^{40},2^{50}\}$, input range from 0 to 2560256 and $N=8$.
We observe that higher values of $\Delta$ yield lower $L_{2}$-norms after a bit flip, indicating stronger attenuation of induced errors. Furthermore, the point at which the error norm begins its exponential growth shifts in alignment with the magnitude of $\Delta$, confirming that larger scaling factors increase the number of error resilient bits in each coefficient.
Figure 10: Histogram of error‐categorization outcomes for different scaling factors. Each group of bars corresponds to a specific $\Delta$, showing the percentage of trials in the different categories. Parameters: 60‑bit $q_{0}$, $\Delta\in\{2^{20},2^{40},2^{50}\}$, input range from 0 to 2560256 and $N=2^{13}$.
In Fig. 10, we adopt a more realistic configuration using a ring dimension $N=2^{13}=8192$ (4096 encryptable elements). For each scaling factor $\Delta\in\{2^{20},2^{40},2^{50}\}$ and each single-bit flip, we track the number of elements that decrypt correctly. We use the three categories defined in Sec. IV-E based on element-wise relative error (Eq. (19)). As $\Delta$ increases, the frequency of Robust outcomes (green bars) also increases, confirming that a larger scaling factor provides greater tolerance to faults.
Although we do not explicitly compute the $L_{2}$-norm threshold at which decryption completely fails, the observed consistency of the Application-dependent error category across scaling factors indicates that the width of the transition interval—from full resilience to total failure—remains approximately constant for all values of $\Delta$. Within this transition region, some elements decrypt correctly while others fail, reflecting the gradual loss of robustness rather than an abrupt threshold.
We observe a slight increase in the Application-dependent category for $\Delta=2^{20}$ and more markedly for smaller $\Delta$ (data not shown). This is consistent with the chosen parameters: with $q_{0}=$60\text{\,}\mathrm{bit}$\$ and $\Delta=2^{20}$ the scheme provides roughly 20 fractional bits and about 40 integer bits. When input values are limited to the range up to $2^{8}=256$, most of the available integer range remains unused, so small perturbations introduced during encoding/decoding (rounding, or bit-flip noise) can move some elements to the Application-dependent category.
V-A3 Impact of the Input Value’s Magnitude
We fix $\Delta=2^{20}$ and $q_{0}=$60\text{\,}\mathrm{bit}$\$ to enable a wider representable input range. As previously explained, the number of bits available to represent the integer part of an input value is approximately $q_{0}-\log_{2}(\Delta)$–about 40 bits in this configuration. In our experiments, we define input value intervals, from which we uniformly sample the elements.
Figure 11 shows histograms of the error category distributions for three input intervals: $2^{9},2^{10}]$, $2^{19},2^{20}]$, and $2^{29},2^{30}]$. We deliberately avoid the interval $2^{39},2^{40}]$ because it may overflow and cause a failed decryption even with unmodified inputs (i.e., without bit flips). The selected intervals are representative of the range of input magnitudes considered in our experiments. We use the following parameters: ring dimension $N=8192$, ciphertext modulus $q_{0}=$60\text{\,}\mathrm{bit}$\$, and scaling factor $\Delta=2^{20}$.
Figure 11: Histogram of error‐categorization outcomes for different scaling inputs. Each group of bars corresponds to a specific input range, showing the percentage of trials in the different categories. Parameters: 60‑bit $q_{0}$, $\Delta=2^{20}$, input range $\in[2^{9},2^{10}],[2^{19},2^{20}]$ and $2^{29},2^{30}]$ and $N=2^{13}$.
@@ V-A3 Impact of the Input Value’s Magnitude
V-A4 Effect of Reducing the Number of Encrypted Elements
As noted previously, CKKS can encode up to $N/2$ elements, where $N$ is the power-of-two ring degree. When we encode fewer elements–still a power of two–we observe that additional coefficients remain unaffected by single‐bit flips. Figure 12 shows results for $N=16$ with only $N/4=4$ encrypted elements.
Figure 12: Bit flip during encryption on vanilla mode. The horizontal axis indicates the bit position in the ciphertext, while the vertical axis shows the resulting $L_{2}$-norm of the decoding error. Parameters: 60‑bit $q_{0}$, $\Delta=2^{20}$, sample input range from 0 to 2560256, $N=16$, and four encrypted elements.
The results reveal that odd‐indexed coefficients of polynomial $c_{0}$ remain unchanged under bit flips, whereas even‐indexed coefficients exhibit the same sensitivity patterns observed earlier.
As explained in Section II, the ratio between the maximum number of encryptable elements ($N/2$) and the number actually used defines an interval (or gap) between the coefficients participating in the inverse transform. During the special FFT, only coefficients which indices are multiple of this gap are considered and scaled in the input vector (see Section I for details). For example, when encoding $N/2$ elements, all $N$ coefficients of $c_{0}$ are sensitive to bit flips, but encoding $N/4$ elements reduces the number of sensitive coefficients of $c_{0}$ to $N/2$. In the case of $c_{1}$, all coefficients are sensitive to bit changes.
V-B Resilience in RNS-only mode
We focus on the RNS-only configuration when the number of encrypted elements is reduced to half or less of the available encryptable elements. Figure 13 reports results for ring dimension $N=16$ with only four encrypted elements, scaling factor $\Delta=2^{20}$, and 8‐bit input values.
Figure 13: Bit flip during encryption on RNS-only mode with two limbs. The horizontal axis indicates the flipped bit position; the vertical axis indicates whether decoding succeeded (robust category) or failed (catastrophic category). Parameters: 60‑bit $q_{0}$, input range from 0 to 2560256, $\Delta=2^{20}$, $N=16$, and four encrypted elements.
We observe that odd‐indexed coefficients of $c_{0}$ remain fully robust: all the input elements decrypt correctly. However, any bit flip in an even‐indexed coefficient, causes catastrophic failures. The number of sensitive coefficients exactly matches the gap analysis from Section V-A4. Polynomial $c_{1}$ exhibits no robust bits at all: every bit flip causes decrypt failure.
These results show that running the RNS-only mode leaving some unused elements can unveil robustness. To ensure a secure-enough schemes, CKKS requires a ring dimension in the order of $2^{15}$, which leads to a payload of near $2^{14}$ encrytable elements. In many practical uses, the amount of needed elements is significantly less, allowing a fault tolerance scheme by taking advantage of the unsed input elements.
MNIST Example. To illustrate the practical impact of bit-flip resilience in CKKS, we consider a case using input images from the MNIST dataset [23]. Each image consists of $28\times 28=784$ grayscale pixels, where each pixel is encoded using integer values ranging from 0 to 255.
To encode 784 values using CKKS batching, we need a vector size that is a power of two and greater than or equal to 784; the smallest size is $2^{10}=1024$. Since CKKS requires a ring dimension $N$ that is twice the batch size, we need $N\geq 2^{11}=2048$. However, to comply with security requirements, we select a ring dimension of $N=2^{15}$. Only $2^{11}$ coefficients of the plaintext or ciphertext polynomials store actual image data, while the remaining $2^{15}-2^{11}$ coefficients are unused.
Figure 14 shows results for $N=2^{15}$, and $2^{10}$ encrypted elements of MNIST images, $q_{0}$ modulus of 60 bits, scaling factor of $2^{20}$, and three-limb RNS. As expected, flipping any bit in the $c_{1}$ polynomial consistently produces an undecryptable result, whereas in $c_{0}$ only about one in every sixteen coefficients is sensitive to bit-flips.
@@ V-B Resilience in RNS-only mode
Figure 14: Bit flip during encryption on RNS-only mode with three limbs, using MNIST images as input. The horizontal axis indicates the flipped bit position; the vertical axis indicates the resulting L2 norm of the decoding error. Only $c_{0}$ is shown because $c_{1}$ is all undecryptable similarly as Fig. 13. Parameters: 60-bit $q_{0}$, $\Delta=2^{20}$, $N=2^{15}$, and $2^{10}$ encrypted elements.
However, this resiliency comes at a substantial computational cost. As explained in Section II, NTT-based multiplication enables fast polynomial product computations with a complexity of $\mathcal{O}(N\log N)$, where $N$ is the ring dimension. When NTT is disabled, the underlying polynomial multiplication reverts to a schoolbook approach with complexity $\mathcal{O}(N^{2})$. As a result, homomorphic operations become orders of magnitude slower when operating in coefficient space without NTT.
Although we do not perform direct runtime benchmarking in this study, prior works such as [6] report performance slowdowns of approximately ten times when switching from the RNS+NTT-optimized scheme to the vanilla CKKS implementation. In scenarios where RNS is used but NTT is disabled, the degradation is expected to be smaller, since RNS still reduces the cost of large-integer arithmetic by enabling parallel residue computations. In practice, the slowdown depends on various factors, including the specific implementation, the chosen ring dimension, and the availability of hardware-level optimizations (e.g., SIMD instructions or parallel FFTs). This highlights a fundamental trade-off between fault resilience and computational efficiency, which must be carefully balanced in real-world deployments.